# LLM4Cell: A Survey of Large Language and Agentic Models for Single-Cell Biology

### v2

@@ Appendix E Methods Comparison
EpiBERT Javed et al. (2025) Cell Genomics Epigenomic Foundation Model / Multi-modal Transformer DNA sequence + chromatin accessibility Masked-accessibility pretraining + motif & sequence fusion No Accessibility imputation, gene expression prediction & regulatory inference - predicts masked ATAC signals, then fine-tunes to predict expression and enhancer-gene links, generalizing to unseen cell types 8
@@
scMamba Yuan et al. (2025) ArXiv Multimodal / Foundation_FM Multi-omics (RNA + others) Implicit via integrated features; no explicit external grounding No Multi-omic embedding, integration & annotation across modalities without prior feature selection 7
@@
GeneMamba Qi et al. (2025a) ArXiv Foundation Model / State-Space Model RNA (scRNA) Implicit (via gene sequence context + pathway loss) No Multi-batch integration, cell-type annotation, gene correlation - uses a BiMamba state-space architecture with pathway-aware losses for scalable, context-rich modeling of scRNA 7
@@
Nicheformer Schaar et al. (2024) BioRxiv Foundation _FM/ Spatial / Multimodal_FM scRNA + spatial transcriptomics Contextual tokenization + metadata + spatial neighborhood embedding No Spatial context prediction, spatial label / niche prediction, mapping spatial info to dissociated cells 7
@@
scFormer Cell+ Cui et al. (2022) Bioarxiv Foundation Model / Multimodal Transformer RNA (+ metadata) Joint gene-cell embedding with metadata tokens No Integration & Annotation - context-aware joint embedding for cross-species/tissue generalization 7
@@
scPlantLLM Cao et al. (2025) Genomics, Proteomics and Bioinformatics Text-Bridge LLM (Foundation Model) RNA (plant scRNA-seq) Gene token + binned expression embedding (Plant Atlas grounding) No Annotation, Integration, GRN inference - pretrains on plant single-cell data with masked LM and cell-type supervision, enabling cross-species annotation, clustering, and regulatory discovery in plant systems 7
@@
LICT Ye et al. (2024) ArXiv Text-Bridge LLM / Annotation LLM Hybrid RNA (scRNA-seq) Marker gene–based DE lists + LLM prompting No Annotation & Reliability - iterative LLM prompting with DE markers for label refinement and confidence scoring 5
@@
Teddy (family of models) Chevalier et al. (2025) ArXiv Foundation Model / Disease-aware Transformer scRNA (single-cell RNA-seq) Self-supervised + supervised annotation supervision No Disease state classification / healthy vs diseased detection - trained to identify disease conditions of held-out donors and distinguish diseased vs healthy cells in new disease contexts 7
@@
Pilot Joodaki et al. (2024) Github Benchmark / Evaluation Framework RNA Model-agnostic pilot foundation testing No Benchmarking & Evaluation - lightweight pilot framework for early single-cell FM testing 4
@@
CellVerse Zhang et al. (2025b) ArXiv Benchmark / Evaluation Multi-omics (RNA, CITE, ASAP, etc.) Implicit via prompt encoding No QA benchmark for annotation, drug response, perturbation tasks 5
@@
xTrimoGene Gong et al. (2023) NeurIPS / ArXiv Foundation Model (Scalable Transformer) scRNA-seq Sparse masking + auto-discretization No Representation learning + annotation, perturbation, drug synergy prediction 7
@@
EpiAttend Li et al. (2022) NeurIPS 2022 Workshop Regulatory / Sequence-Epigenome Transformer DNA sequence + single-cell epigenomic data Sequence + cell-specific epigenome grounding No Predict cell type–specific gene expression by integrating DNA sequence and single-cell epigenomic tracks, linking enhancers and promoters 6
@@
Spatial2-Sentence Chen et al. (2025a) ArXiv Spatial / Text-Bridge hybrid Spatial + expression (Imaging Mass Cytometry) Spatial adjacency + expression similarity tokenization No Encode spatial & expression context into multi-sentence prompts for LLMs to perform cell-type classification and clinical status prediction 6
@@
ChatCell Fang et al. (2024) ArXiv Text-Bridge LLM (Instructional) scRNA → “cell sentence” Vocabulary adaptation + unified sequence generation of cell sentences No Natural language interface for single-cell tasks - allows users to query, annotate, generate, and explore scRNA data via text prompts. Hugging Face 6
@@
CellAtria Nouri et al. (2025) BioRxiv Agentic Framework scRNA- seq + metadata Ontology (graph + metadata) Yes Annotation & Ontology Mapping 7
@@
CellAgent Xiao et al. (2024) ArXiv Agentic Framework scRNA- seq None Yes Annotation & Ontology Mapping 7

### v3

@@ Appendix E Methods Comparison
EpiBERT javed2025multi Cell Genomics Epigenomic Foundation Model / Multi-modal Transformer DNA sequence + chromatin accessibility Masked-accessibility pretraining + motif & sequence fusion No Accessibility imputation, gene expression prediction & regulatory inference - predicts masked ATAC signals, then fine-tunes to predict expression and enhancer-gene links, generalizing to unseen cell types 8
@@
scMamba yuan2025scmamba ArXiv Multimodal / Foundation_FM Multi-omics (RNA + others) Implicit via integrated features; no explicit external grounding No Multi-omic embedding, integration & annotation across modalities without prior feature selection 7
@@
GeneMamba qi2025bidirectional ArXiv Foundation Model / State-Space Model RNA (scRNA) Implicit (via gene sequence context + pathway loss) No Multi-batch integration, cell-type annotation, gene correlation - uses a BiMamba state-space architecture with pathway-aware losses for scalable, context-rich modeling of scRNA 7
@@
Nicheformer schaar2024nicheformer BioRxiv Foundation _FM/ Spatial / Multimodal_FM scRNA + spatial transcriptomics Contextual tokenization + metadata + spatial neighborhood embedding No Spatial context prediction, spatial label / niche prediction, mapping spatial info to dissociated cells 7
@@
scFormer Cell+ cui2022scformer Bioarxiv Foundation Model / Multimodal Transformer RNA (+ metadata) Joint gene-cell embedding with metadata tokens No Integration & Annotation - context-aware joint embedding for cross-species/tissue generalization 7
@@
scPlantLLM scplantlmm Genomics, Proteomics and Bioinformatics Text-Bridge LLM (Foundation Model) RNA (plant scRNA-seq) Gene token + binned expression embedding (Plant Atlas grounding) No Annotation, Integration, GRN inference - pretrains on plant single-cell data with masked LM and cell-type supervision, enabling cross-species annotation, clustering, and regulatory discovery in plant systems 7
@@
LICT ye2024objectively ArXiv Text-Bridge LLM / Annotation LLM Hybrid RNA (scRNA-seq) Marker gene–based DE lists + LLM prompting No Annotation & Reliability - iterative LLM prompting with DE markers for label refinement and confidence scoring 5
@@
Teddy (family of models) chevalier2025teddy ArXiv Foundation Model / Disease-aware Transformer scRNA (single-cell RNA-seq) Self-supervised + supervised annotation supervision No Disease state classification / healthy vs diseased detection - trained to identify disease conditions of held-out donors and distinguish diseased vs healthy cells in new disease contexts 7
@@
Pilot joodaki2024detection Github Benchmark / Evaluation Framework RNA Model-agnostic pilot foundation testing No Benchmarking & Evaluation - lightweight pilot framework for early single-cell FM testing 4
@@
CellVerse zhang2025cellverse ArXiv Benchmark / Evaluation Multi-omics (RNA, CITE, ASAP, etc.) Implicit via prompt encoding No QA benchmark for annotation, drug response, perturbation tasks 5
@@
xTrimoGene gong2023xtrimogene NeurIPS / ArXiv Foundation Model (Scalable Transformer) scRNA-seq Sparse masking + auto-discretization No Representation learning + annotation, perturbation, drug synergy prediction 7
@@
EpiAttend li2022epiattend NeurIPS 2022 Workshop Regulatory / Sequence-Epigenome Transformer DNA sequence + single-cell epigenomic data Sequence + cell-specific epigenome grounding No Predict cell type–specific gene expression by integrating DNA sequence and single-cell epigenomic tracks, linking enhancers and promoters 6
@@
Spatial2-Sentence chen2025spatial ArXiv Spatial / Text-Bridge hybrid Spatial + expression (Imaging Mass Cytometry) Spatial adjacency + expression similarity tokenization No Encode spatial & expression context into multi-sentence prompts for LLMs to perform cell-type classification and clinical status prediction 6
@@
ChatCell fang2024chatcell ArXiv Text-Bridge LLM (Instructional) scRNA → “cell sentence” Vocabulary adaptation + unified sequence generation of cell sentences No Natural language interface for single-cell tasks - allows users to query, annotate, generate, and explore scRNA data via text prompts. Hugging Face 6
@@
CellAtria nouri2025agentic BioRxiv Agentic Framework scRNA- seq + metadata Ontology (graph + metadata) Yes Annotation & Ontology Mapping 7
@@
CellAgent xiao2024cellagent ArXiv Agentic Framework scRNA- seq None Yes Annotation & Ontology Mapping 7